OpenStudy (anonymous):

y= 5/ x+14 solve for y.

OpenStudy (anonymous):

I believe you already have.

OpenStudy (anonymous):

I'm trying to find the inverse of h(x)= 5/x+14 so setting it equal to y and solving for y I'm having troubles with

OpenStudy (anonymous):

\[h ^{-1}(y) = 5/ x+14\]

OpenStudy (anonymous):

i think you want to write \[x=\frac{5}{y+14}\] and solve for \(y\) to find the inverse

OpenStudy (anonymous):

in the equation \(y=\frac{5}{x+14}\) the \(y\) is already solved for, since it is by itself on one side of the equal sign

OpenStudy (anonymous):

do you know how to solve \(x=\frac{5}{y+14}\) for \(y\) ?

OpenStudy (anonymous):

All the answers I am getting from calculating it out seem to be incorrect, maybe step by step would help to see what im doing wrong